# An overview of maximal distance minimizers problem

### v2

@@ 4.3 Convexity argument
Suppose that we fix some $M_{0}\subset M$ and consider a (possibly infinite) tree $T$ whose vertices are encoded by points of $M_{0}$. Let us pick an arbitrary point from $\overline{B_{r}(m)}$ for every $m\in M_{0}$ and connect such points by segments with respect to $T$. Consider the length $L$ of such a representation of $T$; note that we allow the representation to contain cycles or edges of zero length.
@@ 4.5 Results on Steiner trees
We call a topology $T$ realizable for a set $P$ if there exists such a locally minimal tree $S(P)$ with topology $T$; we will denote this tree by $S_{T}(P)$; by the following proposition the notation is correct.
@@ 5.1 $\Gamma$-convergence
$\Gamma$-convergence is an important tool in studying minimizers based on approximation of energy. For Euclidean space the following definition of $\Gamma$-convergence can be used. Let $X$ be a first-countable space and $F_{n}$: $X\to\overline{\mathbb{R}}$ be a sequence of functionals on $X$. Then $F_{n}$ are said to $\Gamma$-converge to a $\Gamma$-limit $F$ : $X\to\overline{\mathbb{R}}$ if the following two conditions hold.
@@
By the definition of $F_{M}$ and triangle inequalities we have
@@
for every connected $S_{i}$ and $S$. So by (1) every sequence of $S_{i}$ with limit $S$ we have the first condition of $\Gamma$-convergence. For the second condition consider $S_{i}$ being a Steiner tree for a finite $1/i$-network $N_{i}\subset S$. By the definition ${\mathcal{H}}(S_{i})\leq{\mathcal{H}}(S)\leq l$. Again, by (1) $F_{M_{i}}(S_{i})$ converges to $F_{M}(S)$.
@@ 5.2 Approximation by Steiner trees
Unfortunately, in the case of several minimizers for $M$ we cannot be sure that every minimizer is approximated. On the other hand it can be approximated a posteriori. Let $\Sigma$ be a minimizer for $M$ and let $\mathcal{E}_{k}\subset\Sigma$ be a finite $1/k$-network and $\Sigma_{k}$ be an arbitrary solution of the Steiner problem for $\mathcal{E}_{k}$. By the definition we have
@@
Summing up, every maximal distance minimizer is a limit of finite Steiner trees. Similar results are also proved in [1]. More detailed and structural relations of finite Steiner trees and maximal distance minimizer are considered in Section 3.1.
@@ 5.3 NP-hardness
To avoid it Garey, Graham and Johnson [9] introduce a discrete version of the Steiner problem. Of course a minimizer of a new problem does not inherit any geometric properties, in particular we have no $32\pi/$ 3-condition at a branching point. Such a discretization appears to be NP-complete (and so the initial one is NP-hard), namely, Garey, Graham and Johnson used a reduction of the X3C problem to this version of the Steiner problem. The X3C problem is to decide whether a family of 3-sets $\mathcal{F}\subset 2^{[3n]}$ has a subfamily of $n$ sets which covers $[3n]$ (as usual, $[3n]=\{1,2,\dots,3n\}$). It is well-known that X3C is NP-complete.
@@
    if $\mathcal{F}$ has an $n$-set covering then a solution of the Steiner problem for $X(\mathcal{F})$ has the length at most $L$;
@@
    if $\mathcal{F}$ does not have an $n$-set covering then a solution of the Steiner problem for $X(\mathcal{F})$ has the length at least $L+12|X(\mathcal{F})|$.
@@
Now let us repeat Garey–Graham–Johnson rounding in the case of maximal distance minimizers. The following problem is a discrete approximation of Problem 1.2 analogous to the discrete version of Steiner problem used in [9]. Following [9] we replace the length function with its ceiling because it is not known if the problem of determining whether $\sum\sqrt{n_{i}}<L$ is NP or not ($n_{i}$, $L$ are integer).
@@
Let $\mathcal{F}\subset 2^{[3n]}$ be an arbitrary family. Consider the set $X(\mathcal{F})$ from Theorem 5.2. Fix any $r\in\mathbb{N}$ and let $k>10r|X(\mathcal{F})|$ be a large integer number. Define $kX(\mathcal{F})$ as a set homothetic to $X(\mathcal{F})$ with the scale factor $k$. Let $M$ be the set of closest points of $kX(\mathcal{F})$ in the integer grid $\mathbb{Z}^{2}$. Put also $\ell=kL(\mathcal{F})+k$.
@@ 5.4 Penalized form
Clearly every set $T$ which minimizes $\lambda$-penalized problem for some $\lambda$ is a maximal distance minimizer for a given data $M$ and the restriction of energy $r:=F_{M}(T)$. Hence the solutions of this problem inherit all regularity properties of maximal distance minimizers.
@@
Hence it is also make sense to consider local penalized problem: the problem of searching a connected compact set $S$ of a finite length, such that ${\mathcal{H}}^{1}(S)+\lambda F_{M}(S)\leq{\mathcal{H}}^{1}(T)+\lambda F_{M}(T)$ for every connected compact $T$ with $\mathrm{diam}\,(S\triangle T)<\varepsilon$ for sufficiently small $\varepsilon>0$. The solutions of these problems also inherit properties of local maximal distance minimizers.
@@ 5.5 Uniqueness
Note that for every compact $M\subset\mathbb{R}^{d}$ and $r$ equal to the radius of the smallest ball containing $M$, there is a unique point $o$ such that $M\subset\overline{B_{r}(o)}$, id est the solution of Problem 1.2 is unique. For a larger $r$ one has an uncountable number of solutions. This motivate us to consider only small enough $r$. Let us call a finite point configuration $M$ ambiguous if Problem 1.2 has several solutions for $M$ and $r<r_{0}(M)$. The following statement is a straightforward corollary from the main theorem of [2].
@@
Fix $n\geq 4$. The main result of [2] states that the Hausdorff dimension of planar $n$-point configurations with multiply Steiner trees is $12n-$ 1.
Recall that a topology $T$ of a labelled Steiner tree is the corresponding abstract graph; a topology is full if every its terminal has degree 1. We call a topology generic if it has no terminals of degree 3. For a not full generic topology $R$ one can replace vertex $A$ of degree 2 with a Steiner point $b$ and add edge $bA$; the resulting topology $T(R)$ is full.
@@
Let us show that if a Steiner tree for a finite $M$ is unique then $M$ is not ambiguous. Consider any $n$-point planar configuration $M$ with unique Steiner tree ${\mathcal{S}t}$ whose topology is generic. Let the length of the second locally minimal tree be ${\mathcal{H}}^{1}({\mathcal{S}t})+a$ and choose $r<a/(2n)$.
@@ 6.1 Regularity
The first question, especially if the answer in negative, might be difficult.
@@
Does there exist a nonplanar maximal distance minimizer with infinite number of branching points?
@@
Question 6.3.
Does Lemma 2.5 holds for a $d$-dimensional maximal distance minimizer?
All the questions in this subsection can be also asked for a local minimizers.
@@ 6.2 Explicit solutions
Define a stadium to be the boundary of the $R$-neighborhood of a segment. By the definition, a stadium has the minimal radius of curvature $R$. Let us show that if $R<1.75r$ and a stadium is long enough, then there is the connected set $\Sigma^{\prime}$ that has the length smaller than an arbitrary horseshoe and covers $M$.
@@
Analogously to the stadium case one can easily show that for some sufficiently small $\frac{|a_{1}a_{2}|}{|a_{2}a_{3}|}<1$ and some $r>0$ a minimizer should have another topology than depicted at Fig. 7.
Also one may consider the following relaxation of Problem 6.6.
Question 6.7.
@@ 6.3 Uniqueness
Conflict of interests.
On behalf of all authors, the corresponding author states that there is no conflict of interest.

### v3

@@ 4.3 Convexity argument
Suppose that we fix some $M_{0}\subset M\subset\mathbb{R}^{d}$ and consider a (possibly infinite) abstract tree $T$ whose vertices are encoded by points of $M_{0}$. Let us pick an arbitrary point from $\overline{B_{r}(m)}$ for every $m\in M_{0}$ and connect such points by segments with respect to $T$. Consider the length $L$ of such a representation of $T$; note that we allow the representation to contain cycles or edges of zero length.
@@ 4.5 Results on Steiner trees
We call a topology $T$ realizable for a set $P\subset\mathbb{R}^{d}$ if there exists such a locally minimal tree $S(P)$ with topology $T$; we will denote this tree by $S_{T}(P)$; by the following proposition the notation is correct.
@@ 5.1 $\Gamma$-convergence
$\Gamma$-convergence is an important tool in studying minimizers based on approximation of energy. For Euclidean space the following definition of $\Gamma$-convergence can be used. Let $X$ be a first-countable space (every Euclidean space is first-countable) and $F_{n}$: $X\to\overline{\mathbb{R}}$ be a sequence of functionals on $X$. Then $F_{n}$ are said to $\Gamma$-converge to a $\Gamma$-limit $F$ : $X\to\overline{\mathbb{R}}$ if the following two conditions hold.
@@
By the definition of $F_{M}$ and the triangle inequality we have
@@
for every connected $S_{i}$ and $S$. So by (1) for every sequence of $S_{i}$ with limit $S$ we have the first condition of $\Gamma$-convergence holds. For the second condition consider $S_{i}$ being a Steiner tree for a finite $1/i$-network $N_{i}\subset S$. By the definition ${\mathcal{H}}(S_{i})\leq{\mathcal{H}}(S)\leq l$. Again, by (1) $F_{M_{i}}(S_{i})$ converges to $F_{M}(S)$.
@@ 5.2 Approximation by Steiner trees
Unfortunately, in the case of several minimizers for $M$ we cannot be sure that every minimizer is approximated. On the other hand it can be approximated a posteriori. Let $\Sigma$ be a minimizer for $M$ and let $\mathcal{E}_{k}\subset\Sigma$ be a finite $1/k$-network and $\Sigma_{k}$ be an arbitrary solution to the Steiner problem for $\mathcal{E}_{k}$. By the definition we have
@@
Summing up, every maximal distance minimizer is a limit of finite Steiner trees. Similar results are also proved in [1]. Recall that more detailed and structural relations of finite Steiner trees and maximal distance minimizers are considered in Section 3.1.
@@ 5.3 NP-hardness
To avoid it Garey, Graham and Johnson [9] introduce a discrete version of the Steiner problem: all terminals and branching points are forced to have integer coordinates and the length of every segment is replaced with its ceiling. Of course a minimizer of a new problem does not inherit any geometric properties, in particular we have no $32\pi/$ 3-condition at a branching point. Such a discretization appears to be NP-complete (and so the initial one is NP-hard), namely, Garey, Graham and Johnson used a reduction of the X3C problem to this version of the Steiner problem. The X3C problem is to decide whether a family of 3-sets $\mathcal{F}\subset 2^{[3n]}$ has a subfamily of $n$ sets which covers $[3n]$ (as usual, $[3n]=\{1,2,\dots,3n\}$). It is well-known that X3C is NP-complete.
@@
    if $\mathcal{F}$ has an $n$-set covering then a solution to the Steiner problem for $X(\mathcal{F})$ has the length at most $L$;
@@
    if $\mathcal{F}$ does not have an $n$-set covering then a solution to the Steiner problem for $X(\mathcal{F})$ has the length at least $L+12|X(\mathcal{F})|$.
@@
Now let us repeat Garey–Graham–Johnson rounding in the case of maximal distance minimizers. The following problem is a discrete approximation of Problem 1.2 analogous to the discrete version of Steiner problem used in [9]. Following [9] we replace the length function with its ceiling because it is not known if the problem of determining whether $\sum\sqrt{n_{i}}<L$ is NP or not ($n_{i}$, $L$ are given integers).
@@
Let $\mathcal{F}\subset 2^{[3n]}$ be an arbitrary family. Consider the set $X(\mathcal{F})$ from Theorem 5.2. Fix any $r\in\mathbb{N}$ and let $k>10r|X(\mathcal{F})|$ be a large integer number. Define $kX(\mathcal{F})$ as a set homothetic to $X(\mathcal{F})$ with the scale factor $k$. Let $M$ be the set of points closest to $kX(\mathcal{F})$ in the integer grid $\mathbb{Z}^{2}$. Put also $\ell=kL(\mathcal{F})+k$.
@@ 5.4 Penalized form
Clearly every set $T$ which minimizes $\lambda$-penalized problem for some $\lambda$ is a maximal distance minimizer for a given input $M$ and the restriction of energy $r:=F_{M}(T)$. Hence the solutions to this problem inherit all regularity properties of maximal distance minimizers.
@@
Hence it is also make sense to consider local penalized problem: the problem of searching a connected compact set $S$ of a finite length, such that ${\mathcal{H}}^{1}(S)+\lambda F_{M}(S)\leq{\mathcal{H}}^{1}(T)+\lambda F_{M}(T)$ for every connected compact $T$ with $\mathrm{diam}\,(S\triangle T)<\varepsilon$ for sufficiently small $\varepsilon>0$. The solutions to these problems also inherit properties of local maximal distance minimizers.
@@ 5.5 Uniqueness
Note that for every compact $M\subset\mathbb{R}^{d}$ and $r$ equal to the radius of the smallest ball containing $M$, there is a unique point $o$ such that $M\subset\overline{B_{r}(o)}$, id est the solution to Problem 1.2 is unique. For a larger $r$ one has an uncountable number of solutions. This motivate us to consider only small enough $r$. Let us call a finite point configuration $M$ ambiguous if Problem 1.2 has several solutions for $M$ and $r<r_{0}(M)$. The following statement is a straightforward corollary from the main theorem of [2].
@@
Fix $n\geq 4$. The main result of [2] states that the Hausdorff dimension of planar $n$-point configurations with multiple Steiner trees is $12n-$ 1.
Recall that a topology $T$ of a labelled Steiner tree is the corresponding abstract graph; a topology is full if every terminal has degree 1. We call a topology generic if it has no terminals of degree 3. For a not full generic topology $R$ one can replace vertex $A$ of degree 2 with a Steiner point $b$ and add edge $bA$; the resulting topology $T(R)$ is full.
@@
Let us show that if a Steiner tree for a finite $M$ is unique then $M$ is not ambiguous. Consider any $n$-point planar configuration $M$ with a unique Steiner tree ${\mathcal{S}t}$ whose topology is generic. Let the length of the second locally minimal tree be ${\mathcal{H}}^{1}({\mathcal{S}t})+a$ and choose $r<a/(2n)$.
@@ 6.1 Regularity
The first question, especially if the answer is negative, might be difficult.
@@
Does there exist a nonplanar maximal distance minimizer with an infinite number of branching points?
The next question concerns the Ahlfors regularity in the following strong form for $d>2$.
Question 6.2.
Are there exist constants $C_{1},C_{2}$ depending only on $d$ such that for every $M$ and $r>0$ there is a $\varepsilon_{0}>0$ for which
$C_{1}\varepsilon\leq{\mathcal{H}}^{1}(\Sigma\cap B_{\varepsilon}(x))\leq C_{2}\varepsilon$
holds for every $x\in\Sigma$ and $\varepsilon<\varepsilon_{0}$?
@@
Question 6.4.
Does Lemma 2.5 hold for a $d$-dimensional maximal distance minimizer?
All the questions in this subsection can be also asked for local minimizers.
@@ 6.2 Explicit solutions
Define a stadium to be the boundary of the $R$-neighborhood of a segment. By the definition, a stadium has the minimal radius of curvature $R$. Let us show that if $R<1.75r$ and a stadium is long enough, then there is the connected set $\Sigma^{\prime}$ that covers $M$ and has the length smaller than an arbitrary horseshoe.
@@
Analogously to the stadium case one can easily show that for a narrow rectangle (i.e. with a sufficiently small $\frac{|a_{1}a_{2}|}{|a_{2}a_{3}|}$) and some $r>0$ a minimizer should have another topology than depicted at Fig. 7.
Also one may consider the following relaxation of Problem 6.7.
Question 6.8.
@@ 6.3 Uniqueness
Recall that we believe that every Steiner tree which is not unique for its $n$ terminals cannot be a maximal distance minimizer for any $n$-point set $M$ and any $r>0$. This motivated the following more general question.
Question 6.11.
Fix $M\subset\mathbb{R}^{d}$ and $r>0$. Does a set $G_{\Sigma}$ of energetic points determine a minimizer $\Sigma$?